# COBRA-PPM: A Causal Bayesian Reasoning Architecture Using Probabilistic Programming for Robot Manipulation Under Uncertainty

### 24 Jun 2025 v3

@@ IV-A2 Unlocking the Power of the Python Ecosystem
We implement the robot decision-making model as a generative probabilistic program in Pyro, a probabilistic programming language (PPL) built on Python. This enables integration with a range of scientific and robotics packages, allowing the construction of causal models using existing libraries.
In our exemplar task, the model leverages the PyBullet simulator during inference-time sampling to evaluate candidate actions and task outcomes. Beyond physics, Python provides access to scientific libraries for tasks such as signal modelling (e.g., for lidar, image, or wireless propagation), object classification, human pose estimation, and cognitive models for human-robot interaction (e.g., theory of mind or intent estimation). These components can be embedded directly into the causal model to support reasoning over perceptual, social, and contextual variables.
@@ IV-B Intervention-Based Causal Inference
A second key component of our architecture is intervention-based causal inference, used to answer prediction and action-selection queries. Representing the model in Pyro enables both exact and approximate inference methods, including sample-based techniques such as importance sampling [26] and gradient-based approaches like stochastic variational inference (SVI) [27]. Pyro also permits flexible composition of conditioning statements with interventions via the $do(\cdot)$ operator, allowing estimation of interventional transition posteriors conditioned on robot observations. These posteriors support predictive queries over current and future task states for decision-making.
@@ V-B2 Task State Representation
The task state at time $k$, denoted $s_{k}$, consists of the current tower configuration $t_{k}$ and the queue of remaining blocks $q_{k}$: $s_{k}=t_{k}\cup q_{k}$. An example showing the initial state $s_{0}$ and resulting state $s_{1}$ after action $a_{1}$ is illustrated in Fig. 3. The tower state $t_{k}$ is represented as an ordered list of block references: $t_{k}=[b_{n_{i},k}]$ for $i=0,\dots,I_{k}$, where $n_{i}$ is the unique identifier of the block at position $i$ in the tower, and $I_{k}+1$ is the number of blocks in the tower at time $k$. Similarly, the queue state is: $q_{k}=[b_{n_{j},k}]$ for $j=0,\dots,J_{k}$, where $n_{j}$ is the unique identifier of the block at position $j$ in the queue, and $J_{k}+1$ is the number of remaining blocks in the queue. Each block state $b_{n,k}$ encodes the physical attributes of block $n$ at time $k$, including its 6-DOF pose (position and orientation), 3D dimensions, and mass.
@@ V-B4 Transition Function Representation
We model the state transition function $T$ following the standard POMDP formulation [23], adopting a causal perspective by treating the agent’s action as an intervention using the $do(\cdot)$ operator. This allows $A_{k}$ to be interpreted as a direct manipulation of the system, marginalising over execution noise $W_{A,k}$: $T(S_{k-1},A_{k},S_{k})=P\big(S_{k}\mid S_{k-1},do(A_{k}=a_{k})\big)$. To sample the successor state $s_{k}$ from the transition distribution $T$, we use the PyBullet 3D physics simulator [28] to simulate the 6-DOF rigid body dynamics of the tower (Fig. 4), conditioned on the agent’s noisy observation $Z_{k-1}$ of the previous state.
@@ V-B5 Noise & Error Modelling
We model both observation and manipulation errors as zero-mean, isotropic Gaussian noise in 3D. Specifically, observation noise and action execution error are represented as independent random vectors with components drawn from $\mathcal{N}(0,\sigma_{Z}^{2})$ and $\mathcal{N}(0,\sigma_{A}^{2})$, respectively. The standard deviations $\sigma_{Z}$ and $\sigma_{A}$ represent uncertainty in the robot’s perception and control, due to sensor noise, estimation errors, and actuation imprecision. All noise components are assumed independent across the $x$-, $y$-, and $z$-axes and over time.
@@ VI-A High-Fidelity Gazebo Robot Simulation Evaluation
We evaluate our architecture on the exemplar block stacking task using a simulated Toyota Human Support Robot (HSR) [29] in Gazebo [30]. The setup uses three 7.5 cm cube blocks matching those in the real-world demonstration (Fig. 2). To simulate realistic block position estimation, we attach ArUco markers to each block and generate 6-DOF pose observations from simulated RGB-D data using the OpenCV ArUco module. Pick-and-place actions are executed via motion plans generated by ROS MoveIt!.
@@ VI-B2 Tower Stability Classification Accuracy
We generate a dataset of 1000 randomly sampled 3-block tower configurations in simulation to evaluate our model’s classification accuracy. Ground-truth stability labels are assigned by forward-simulating each tower configuration in the Gazebo physics engine and observing whether the tower remains standing. We use the Importance Sampling inference method in Pyro, drawing 50 samples per query, to estimate tower stability based on our exemplar task decision-making causal model (see Sec. V-B) and the previously estimated value of $\sigma_{Z}$. Binary classifications are produced by thresholding the predicted stability probability, and evaluated across a range of threshold values. To assess performance, we compute standard classification metrics — $\text{F}_{1}$ score, AUC, precision, and recall — across a range of thresholds.
@@ VII-A1 Block Pose Estimation Error Characterisation
Results from the block position error characterisation are shown in Table I. The data reveal that our initial assumption, that the random error term $W_{Z}$ is independent of the robot’s state, does not strictly hold. In particular, measurement error along the X-axis — approximately aligned with the camera’s depth axis — exhibits a non-zero mean and substantially larger standard deviation than the Y- and Z-axes. Despite this epistemic modelling error, it does not significantly degrade model performance. We therefore retain the simplifying assumption of zero-mean, state-independent noise. Standard deviations along each axis ($\sigma_{Z,x}$, $\sigma_{Z,y}$, $\sigma_{Z,z}$) confirm that the measurement error is approximately Gaussian. We take the average of these values as a sufficiently accurate approximation for a shared standard deviation $\sigma_{Z}=0.469$ cm, applied across all axes in our 3D Gaussian noise model.
TABLE I: Characterisation of block position measurement and placement errors. Values report standard deviation (in cm) of position error along each axis. Model parameters $\sigma_{Z}$ and $\sigma_{A}$ are computed as the mean across axes.
@@ VII-B2 Discussion
The improved performance of our architecture over the baseline likely stems from its ability to explicitly model rigid-body physics and account for uncertainties in perception and manipulation. By integrating the PyBullet 3D physics simulator into the causal model, our method simulates system dynamics with high accuracy, enabling realistic predictions of candidate block placements. In contrast, the baseline uses a simple heuristic that ignores both dynamics and uncertainty.
As a result, our architecture generalises more effectively across varied initial tower configurations. By leveraging mental simulation for online action selection, it remains robust even in challenging cases — e.g., when intermediate blocks are placed off-centre. This robustness is further demonstrated by the perfect 100% task success rate in the ideal setting without manipulation error, highlighting the advantage of our method when precise action selection is critical.
@@ VII-C Architecture Scalability & Limitations
Our architecture is designed to be modular and extensible, enabling its application to a wide range of manipulation scenarios. In this section, we explore its scalability along three axes: action complexity, sequential reasoning, and reward and uncertainty modelling. We also discuss limitations and considerations when scaling to more complex tasks.
@@ VII-C1 Action Spaces
The proposed dynamic CBN-based decision-making causal model can be extended to support richer action representations. In principle, additional action variables — such as diverse block geometries, frictional properties, or non-canonical placement orientations — can be incorporated by expanding the causal graph. Pyro’s plate notation and conditional independence enable efficient factorisation of the joint distribution, improving scalability over flat representations. Our model also supports multi-tower or interacting structure scenarios by modularising local subgraphs and leveraging a physics-based simulator — such as PyBullet, as demonstrated in our exemplar task — to capture the relevant physical interactions.
@@ VII-C3 Reward Functions and Statistical Objectives
The architecture supports flexible objectives beyond expectations over Gaussian rewards. Risk-aware metrics such as CVaR, percentile thresholds, or multi-objective criteria can be incorporated without changes to the underlying causal graph. For example, a practitioner may wish to optimise for both task success and robustness to disturbances such as surface vibrations or wind. Probabilistic programming enables sampling-based estimation under arbitrary distributions, allowing the use of non-Gaussian, asymmetric, or heavy-tailed beliefs. This flexibility supports a wide range of task definitions, robot morphologies, sensing modalities, and deployment environments.
@@ VII-C5 Adaptability and Domain Transfer
Although causal models are interpretable and structured, deploying them in new environments may require domain-specific assumptions or structure learning, introducing some overhead compared to fully data-driven approaches. However, our framework is explicitly designed for modularity, reuse, and abstraction: the causal graph, inference engine, and decision-making components are loosely coupled and can be adapted independently. Further, causal models are well suited to domain transfer, as they aim to capture fundamental causal mechanisms that govern system behaviour and are expected to remain invariant across environments. This composability supports flexible transfer across domains by reusing learned causal templates, swapping simulators, or integrating new inference targets — without redesigning the overall architecture.
@@ VIII Real-World Robot Demonstration
The system succeeded in 4 out of 5 trials (80% success rate), demonstrating strong sim2real generalisation and suitability for robot deployment without re-tuning or retraining.
While initial configurations differ slightly, the core task remains unchanged: reasoning under uncertainty about tower stability and action consequences. One possible source of performance variation is that, in simulation, the robot places a third block onto randomly generated two-block towers, which may be less stable than the deliberately constructed intermediate two-block towers in real-world trials. Nevertheless, our model explicitly accounts for perception and actuation noise, enabling robust performance across sequential placements despite cumulative uncertainty.

### 29 Aug 2025 v4

@@ Abstract
©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ IV-A2 Unlocking the Power of the Python Ecosystem
We implement the robot decision-making model as a generative probabilistic program in Pyro, a probabilistic programming language (PPL) built on Python. Consequently, our PPL-based architecture can handle modelling distributions as arbitrary general-purpose Python code, and is not limited to closed-form analytical functions or differentiable programs as many other probabilistic methods are. In addition to a flexible modelling capability, this enables integration with a range of scientific and robotics packages, allowing the construction of causal models using existing libraries.
In our exemplar task, the model leverages the PyBullet simulator during inference to evaluate candidate actions and task outcomes. Beyond physics, Python provides access to scientific libraries for tasks such as signal modelling (e.g., for lidar, image, or wireless propagation), object classification, human pose estimation, and human cognitive models for human-robot interaction. These components can be embedded directly into the causal model to support reasoning over perceptual, social, and contextual variables.
@@ IV-B Intervention-Based Causal Inference
The modelling flexibility enabled by our PPL-based architecture also extends to inference: once the generative model is defined in Pyro, it can be paired with a range of inference algorithms to support prediction and decision-making over arbitrary-code models. Representing the model in Pyro enables both exact and approximate inference methods, including sample-based (e.g., importance sampling [26]) and gradient-based methods (e.g., stochastic variational inference, SVI [27]). Pyro also permits flexible composition of conditioning statements with interventions via the $do(\cdot)$ operator, allowing estimation of interventional transition posteriors conditioned on robot observations. These posteriors support predictive queries over current and future task states for decision-making. This end-to-end approach enables tractable inference over causal models that incorporate rich, domain-specific simulators and perception modules.
@@ V-B2 Task State Representation
The task state at time $k$, denoted $s_{k}$, consists of the current tower configuration $t_{k}$ and the queue of remaining blocks $q_{k}$: $s_{k}=t_{k}\cup q_{k}$. An example showing the initial state $s_{0}$ and resulting state $s_{1}$ after action $a_{1}$ is shown in Fig. 3. The tower state $t_{k}$ is represented as an ordered list of block references: $t_{k}=[b_{n_{i},k}]$ for $i=0,\dots,I_{k}$, where $n_{i}$ is the unique identifier of the block at position $i$ in the tower, and $I_{k}+1$ is the number of blocks in the tower at time $k$. Similarly, the queue state is: $q_{k}=[b_{n_{j},k}]$ for $j=0,\dots,J_{k}$, where $n_{j}$ is the unique identifier of the block at position $j$ in the queue, and $J_{k}+1$ is the number of remaining blocks. Each block state $b_{n,k}$ encodes the physical attributes of block $n$ at time $k$, including its 6-DOF pose (position and orientation), 3D dimensions, and mass.
@@ V-B4 Transition Function Representation
We model the state transition function $T$ following the standard POMDP formulation [23], adopting a causal perspective by treating the agent’s action as an intervention using the $do(\cdot)$ operator. This allows $A_{k}$ to be interpreted as a direct manipulation of the system, marginalising over execution noise $W_{A,k}$: $T(S_{k-1},A_{k},S_{k})=P\big(S_{k}\mid S_{k-1},do(A_{k}=a_{k})\big)$. Uncertainty in observation and action outcome is introduced by sampling error terms $W_{Z,k}$ and $W_{A,k}$ from their respective noise distributions. To sample the successor state $s_{k}$ from the transition distribution $T$ during CBN-based model sampling and inference, we use the PyBullet 3D physics simulator [28] to simulate the 6-DOF rigid body dynamics of the tower (Fig. 4), conditioned on the agent’s noisy observation $Z_{k-1}$ of the previous state. These samples are propagated into the initial PyBullet simulation state, perturbing the placement pose of the new block away from the robot’s intended action $(x,y)$. The simulator then performs a deterministic forward simulation of the physical system. The resulting tower configuration is used to determine the successor state.
@@ V-B5 Noise & Error Modelling
We model both observation and manipulation errors as zero-mean, isotropic Gaussian noise in 3D. Specifically, observation noise and action execution error are represented as independent random vectors with components drawn from $\mathcal{N}(0,\sigma_{Z}^{2})$ and $\mathcal{N}(0,\sigma_{A}^{2})$, respectively. The standard deviations $\sigma_{Z}$ and $\sigma_{A}$ represent uncertainty in the robot’s perception and control, due to sensor noise, estimation errors, and actuation imprecision. Noise terms are assumed independent across axes and time.
@@ VI-A High-Fidelity Gazebo Robot Simulation Evaluation
We evaluate our architecture on the exemplar block stacking task using a simulated Toyota Human Support Robot (HSR) in Gazebo. The setup uses three 7.5 cm cube blocks matching those in the real-world demonstration (Fig. 2). To simulate realistic block position estimation, we attach ArUco markers to each block and generate 6-DOF pose observations from simulated RGB-D data using the OpenCV ArUco module. Pick-and-place actions are executed via motion plans generated by ROS MoveIt!.
@@ VI-B2 Tower Stability Classification Accuracy
We generate a dataset of 1000 randomly sampled 3-block tower configurations in simulation to evaluate model classification accuracy. Ground-truth stability labels are assigned by forward-simulating each configuration in Gazebo and observing whether the tower remains standing. We use the Importance Sampling inference method in Pyro, drawing 50 samples per query, to estimate tower stability based on our exemplar task causal model (see Sec. V-B) and the estimated value of $\sigma_{Z}$. Binary classifications are produced by thresholding the predicted stability probability. To assess performance, we compute standard classification metrics ($\text{F}_{1}$ score, AUC, precision, and recall) across a range of thresholds.
@@ VII-A1 Block Pose Estimation Error Characterisation
Results from the block position error characterisation are shown in Table I. The data reveal that our initial assumption, that the random error term $W_{Z}$ is independent of the robot’s state, does not strictly hold. In particular, measurement error along the X-axis, approximately aligned with the camera’s depth axis , exhibits a non-zero mean and substantially larger standard deviation than the Y- and Z-axes. Despite this epistemic modelling error, it does not significantly degrade model performance. We therefore retain the simplifying assumption of zero-mean, state-independent noise. The empirical error distributions (not shown due to space constraints) are approximately symmetric and bell-shaped, supporting the use of a Gaussian approximation. We take the average as an accurate estimate for a shared standard deviation $\sigma_{Z}=0.469$ cm, applied across all axes in our 3D Gaussian noise model.
TABLE I: Gazebo-based characterisation of block placement and measurement errors. Values show position error standard deviation (cm) along each axis. Model parameters (in-text) are computed as the mean across axes.
@@ VII-B2 Discussion
The improved performance of our architecture over the baseline likely stems from its ability to explicitly model rigid-body physics and account for uncertainties in perception and manipulation. By integrating the PyBullet 3D physics simulator into the causal model, our method simulates dynamics with high accuracy, enabling realistic predictions of candidate block placements. In contrast, the baseline uses a simple heuristic that ignores both dynamics and uncertainty. Thus, our architecture generalises more effectively across varied initial tower configurations. By leveraging mental simulation for action selection, it remains robust even in challenging cases, e.g., when intermediate blocks are placed off-centre. This robustness is further demonstrated by the 100% success rate in the ideal setting without manipulation error, highlighting the advantage of our method when precise action selection is critical.
@@ VII-C Architecture Scalability & Limitations
Our architecture is designed to be modular and extensible, enabling application to a wide range of manipulation scenarios. In this section, we explore its scalability and discuss limitations when applied to more complex tasks.
@@ VII-C1 Action Spaces
The proposed dynamic CBN-based decision-making causal model can be extended to support richer action representations. In principle, additional action variables, such as diverse block geometries, frictional properties, or non-canonical placement orientations, can be incorporated by expanding the causal graph. Pyro’s plate notation and conditional independence enable efficient factorisation of the joint distribution, improving scalability over flat representations. Our model also supports multi-tower or interacting structure scenarios by modularising local subgraphs and using a physics-based simulator, such as PyBullet (as in our exemplar task), to capture relevant physical interactions.
@@ VII-C3 Reward Functions and Statistical Objectives
The architecture supports flexible objectives beyond expectations over Gaussian rewards. Risk-aware metrics such as CVaR, percentile thresholds, or multi-objective criteria can be incorporated without changes to the causal graph. For example, a practitioner may wish to optimise for both task success and robustness to disturbances such as surface vibrations or wind. Probabilistic programming enables sampling-based estimation under arbitrary distributions, allowing non-Gaussian, asymmetric, or heavy-tailed beliefs. This flexibility supports diverse task definitions, robot morphologies, sensing modalities, and deployment environments.
@@ VII-C5 Adaptability and Domain Transfer
Although causal models are interpretable and structured, deploying them in new environments may require domain-specific assumptions or structure learning, introducing overhead compared to fully data-driven methods. However, our framework is designed for modularity and reuse: the causal graph, inference engine, and decision-making components are loosely coupled and can be adapted independently. Causal models are also well suited to domain transfer, as they aim to capture fundamental causal mechanisms that govern system behaviour and are expected to remain invariant across environments. This composability supports flexible transfer by reusing learned causal templates, swapping simulators, or integrating new inference targets without modifying the overall architecture.
@@ VIII Real-World Robot Demonstration
The system succeeded in 4 out of 5 trials, demonstrating strong sim2real generalisation and suitability for deployment without re-tuning or retraining. While initial configurations differ slightly, the core task remains unchanged: reasoning under uncertainty about tower stability and action consequences. A possible source of variation is that, in simulation, the robot places a third block onto randomly generated two-block towers, which may be less stable than the deliberately constructed intermediate two-block towers in real-world trials. Nevertheless, our model explicitly accounts for perception and actuation noise, enabling robust performance across sequential placements despite cumulative uncertainty.